OpenStudy (anonymous):

What is the sine value of 5 pi over 3?

OpenStudy (help!!!!):

\[\sin \left(\frac{5\pi }{3}\right)=-\frac{\sqrt{3}}{2}\]

OpenStudy (help!!!!):

best response is appreciated

OpenStudy (help!!!!):

If you need help with the steps here is the link

OpenStudy (anonymous):

Can you help me with another?

OpenStudy (help!!!!):

sure just post another question and mention me

OpenStudy (anonymous):

I picked that but I thought it was cos